OpenStudy (theedwardsfamily):

What is the solution to the equation? 2(m-8)=4(m+6) m=

OpenStudy (anonymous):

-20

OpenStudy (theedwardsfamily):

@Albany_Goon

Mehek (mehek14):

2m-16=4m+24 2m=4m+8 -2m=8 m=-4 idk how the other guy got -20